OpenStudy (anonymous):

Suppose you and three friends are equally sharing the total cost C of a meal, and you pay $7. Which equation represents this situation

hartnn (hartnn):

3 friends equally sharing means share of all of them is $7 total cost = 3* ($7) and its given as C so can u find the equation now ?

OpenStudy (anonymous):

ok thx

hartnn (hartnn):

what u got the equation as ?

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

would it be C=4/7 or 7= 4/C

hartnn (hartnn):

ok, there are total 4 people.... so, 4*7 = C is there an option like this ?

OpenStudy (anonymous):

no

hartnn (hartnn):

list all options.

OpenStudy (anonymous):

A C=4/7

OpenStudy (anonymous):

B C= 7/4

OpenStudy (anonymous):

C. C=4/C

OpenStudy (anonymous):

D.7=C/4

hartnn (hartnn):

C=7*4 dividing 4 on both sides 7= c/4

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so 7=C/4

hartnn (hartnn):

yes, option D

OpenStudy (anonymous):

ok thank you